### v2

# High thermal conductivity dominated by thermal phonons with mean free paths exceeding hundred nanometer in three-dimensional covalent organic framework derivatives: a molecular dynamics study

Abstract
Thermal properties of covalent organic frameworks (COFs) are of fundamental interest owing to exceptional heat conduction properties. Recent studies have suggested that their thermal conductivities can be enhanced by multiple factors such as pore size, mass density, and degree of chain order. However, microscopic processes that govern heat conduction properties have been explored in only a limited number of COFs. Here, we report thermal transport properties of 3D COF derivatives using molecular dynamics simulations. In this work, we have studied six different COF-102 derivatives with different organic linkers and topologies. Among the derivatives studied, we found that COF-102 derivatives with high mass density can exhibit thermal conductivity as high as Wm-1K-1, owing to suppressed chain rotation that leads to thermal phonons scattered by anharmonicity. Our results show that the observed orders of magnitude of increase in the thermal conductivity are primarily attributed to low frequency phonon modes that support hundreds of nanometer scale mean free path, which predominantly carry heat. Our study provides a theoretical framework that elucidates the structure-property relationship governing heat conduction in COFs, offering design strategies for thermal management applications.
I Introduction
Covalent organic frameworks (COFs), a class of crystalline porous organic materials, hold promise for various applications such as gas storage and separation Furukawa and Yaghi (2009); Fan et al. (2018), catalysis Wang et al. (2018), optoelectronics Keller and Bein (2021) and energy storage Zhu et al. (2021), owing to their large surface areas, tunable pore sizes, and high thermal stabilities Côté et al. (2005); El-Kaderi et al. (2007); Diercks and Yaghi (2017). Prior extensive studies have suggested that the diverse nature of structure, composition, and functionalities facilitates the tunability of their physical properties for tailored applications Huang et al. (2016); Liu et al. (2021a). For instance, conjugated COFs have shown unique optoelectronic phenomena such as exceptionally high charge carrier mobility Guo et al. (2013) and tunable fluorescence emission as a result of cation addition Ding et al. (2016). Additionally, COFs could exhibit other interesting properties such as facile ion transport Wang et al. (2017); Li et al. (2021) and thermoelectricity Wu et al. (2023); Zhou et al. (2023), which are crucial for energy storage and conversion applications Hu et al. (2020). However, efficient heat dissipation is needed in such applications to ensure device stability Hu and Luo (2019), making it essential to understand heat transport properties in COFs Shakouri (2006); Pop (2010); Babaei et al. (2020).
Thermal conduction in COFs is interesting due to several reasons. First, bulk COFs both in crystalline and amorphous phases exhibit exceptionally low thermal conductivities close to those of many bulk polymers. While lacking existing literature, amorphous COFs, including COF-300 and reticular innovative organic materials (RIO) derivatives, were reported to have thermal conductivities of Wm-1K-1 Freitas et al. (2017) that do not exhibit marked differences with those of crystalline COFs Thakur and Giri (2023); Liu et al. (2016). Second, due to their inherent porosity, COFs generally show a negative correlation between thermal conductivity and pore size, despite few exceptions Ma et al. (2021); Giri and Hopkins (2021). Given the structural tunability of COFs, it has been suggested that their thermal transport properties can be modulated through rational design of linkers and net topologies. Third, constituent light elements (e.g., B, C, N) connected by strong covalent bonds offer possibilities of enhancing heat conduction properties. In fact, prior studies have observed that heat conduction in crystalline porous materials becomes increasingly impeded as linker-node mass mismatch gets larger Islamov et al. (2023). A representative example is metal organic frameworks (MOFs), which unavoidably contain metal nodes that exhibit a large mass mismatch with organic linkers. This leads to additional vibrational scattering, which, in turn, suppresses thermal conductivity Wieser et al. (2021); Zhou et al. (2022). In support of these observations, while MOFs generally exhibit glass-like thermal conductivities typically less than Wm-1K-1 Wang et al. (2015); Zhang et al. (2021); Huang et al. (2019); Babaei et al. (2020), recent experimental work for 2D crystalline COF thin films has reported thermal conductivity exceeding 1 Wm-1K-1 along with high longitudinal speed of sound Evans et al. (2021).
COFs share chemical and structural similarities with polymers, suggesting their potential for enhancing thermal transport properties. According to prior studies, polymers can be thermally conductive if they have a large atomic mass density Xie et al. (2016, 2017), which is associated with densely packed crystalline domains composed of aligned covalent chains Lv et al. (2022). In the extreme, perfect polymer crystals, such as polyethylene (PE) Shulumba et al. (2017) and polythiophene (PT) Cheng et al. (2019), have been reported to exhibit thermal conductivities exceeding Wm-1K-1. Similarly, in COFs, design strategies such as tuning functional groups and pore sizes to increase mass density have been shown to enhance thermal conductivity. Recently, Rahman et al. observed that thermal conductivity of 2D COFs monotonically increases with mass density Rahman et al. (2023). The observed trend was more pronounced in COFs with nodes that have increased connectivity through stronger covalent bonds to the surrounding linkers, resulting in weak anharmonicity. Additionally, Ma et al. demonstrated that COF-300 derivatives can achieve high thermal conductivity of Wm-1K-1 along all spatial directions when pore size is reduced to 0.63 nm Ma et al. (2021). In support of these findings, a recent high-throughput screening study that surveyed over 10,000 3D COFs reported that thermal conductivity as high as Wm-1K-1 can be achieved through enhanced chain orientations and appropriate choice of linkage material Thakur and Giri (2024). These findings suggest that understanding of these design factors could improve the intrinsic upper limit of thermal conductivities in COFs.
Here, we investigate the thermal transport properties of COF-102 derivatives using molecular dynamics (MD) simulations. In this work, we have studied six different COF-102 derivatives with different organic linkers and topologies, which vary elastic stiffness, mass density, and pore size in them. Among the derivatives investigated in this work, we found that COF-102-1 (ctn) that has the highest mass density and speed of sound, can exhibit thermal conductivity as high as Wm-1K-1, nearly two-fold increase compared to the original COF-102, primarily due to suppressed chain rotations, leading to heat-carrying phonons scattered by anharmonicity. The observed increase in the thermal conductivity is primarily due to low frequency phonons with mean free paths (MFPs) on the order of hundreds of nanometers, which are dominant contributors to heat conduction in them. Our study provides a theoretical framework for tailoring thermal properties of 3D COFs for thermal management applications.
II Methods
We first constructed atomic structures of COF-102 derivatives and optimized their energies using the polymer consistent force field (PCFF) Sun et al. (1994) in Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) Plimpton (1995). The detailed structural and physical properties of COF-102 derivatives are provided in Supporting Information Sec. SI. Further MD calculations with periodic boundary conditions (PBC) applied along all spatial directions were conducted using PCFF potential in LAMMPS.
| Material | Topology | Number of unit cells | Supercell size (nm) |
| COF-102-1 | ctn | 4.45 | |
| bor | 4.27 | ||
| COF-102-2 | ctn | 5.44 | |
| bor | 5.28 | ||
| COF-102-3 | ctn | 8.59 | |
| bor | 8.43 |
For EMD simulations, the thermal conductivity was computed using the Green-Kubo formalism Chen (2005).
Here, is system volume, is the Boltzmann constant, is temperature, and is heat current. The term represents ensemble average of heat current autocorrelation function (HCACF) that decays with time. Thermal conductivity was calculated by integrating HCACF over a correlation time that is sufficiently longer than time at which HCACF decays to zero (see Supporting Information Sec. SII for details).
Simulated systems were first relaxed in the NPT, NVT, then NVE ensemble at 300 K for 1 ns. Then, heat current data were collected in the additional NVE ensemble for 5 ns. To reduce statistical errors, thermal conductivities were averaged over eight independent simulation runs that started with different initial velocities. In all EMD simulations, a time step of 1 fs was used. The sizes of simulated structures are summarized in Tab. 1.
Phonon dispersion relations were calculated by using SED Thomas et al. (2010) as a function of angular frequency and wave vector , which can be expressed as
where is total integration time, is total number of unit cells, and are mass and velocity of atom with index , and is equilibrium position of each unit cell.
In our SED analysis, we used systems with , in which corresponds to the number of unit cells used in EMD along one spatial direction; the systems underwent identical relaxation process employed in EMD simulations at 300 K, except the additional subsequent NVE ensemble at which velocities and positions of atoms were recorded for 4 ns to calculate the SED. The resulting SED values were averaged over two independent runs. Phonon properties were computed by fitting frequency spread for each phonon mode to Lorentzian function
where is peak intensity, is frequency at peak position, and is half-width at half-maximum. As established in the literature Thomas et al. (2010), MFPs of phonon modes were determined by , where phonon lifetime is given by and phonon group velocity is defined as .
In NEMD simulations, we varied domain size along the axial direction () from 33.1 to 198.3 nm, while keeping lateral dimensions equivalent to those used in EMD simulations. The thermal conductivity was computed by using Fourier’s law Liu et al. (2021b); Chakraborty et al. (2020):
where is heat flux and is temperature gradient. System was first relaxed in the NPT ensemble at 300 K for 1 ns. Subsequently, two ends with a length of 2.2 nm were set as fixed regions. Langevin thermostats were applied to heat source and sink regions (length: 8.8 nm) to generate steady-state heat flux along the axial direction under a temperature gradient for 4 ns; the data (temperature gradient and heat flux) obtained after initial 1 ns were used to compute the thermal conductivity. Temperatures of heat source and sink regions placed next to each fixed region were set to be at 330 K and 270 K, respectively. A time step of 0.25 fs was used in our entire NEMD calculations.
III Results and discussion
We studied the thermal transport of COF-102 derivatives using MD simulations. Detailed descriptions of the MD simulations implemented in this work are given in Supporting Information Sec. SII and II Methods. As illustrated in Fig. 1(a), the COF-102 derivatives studied in this work were constructed from two cubic topologies, carbon nitride (ctn) and boracite (bor). They are formed by linking tetrahedral carbon nodes with 3-connected linkers; these nodes are connected to three distinct linkers via C-C bonds. The size of linker depends on the number of phenyl rings it contains. Although both ctn and bor topologies share the same stoichiometry, bor topologies are generally % less dense than ctn topologies (see Supporting Information Sec. SI for details).
Figure 1(b) shows room temperature thermal conductivities of COF-102 derivatives, calculated using the Green-Kubo formalism (Eq. 1). As expected, a positive correlation between thermal conductivity and mass density is clearly visible. We note that the opposite trend is observed between thermal conductivity and pore size (see Supporting Information Fig. S3 in Sec. SII). COF-102-1, in particular, exhibits the highest thermal conductivity ( Wm-1K-1; Wm-1K-1), while the others show thermal conductivities reduced by nearly two orders of magnitude. In all cases, bor topologies yield % lower thermal conductivities than ctn topologies. We here note that COF-102-2 (ctn) is structurally similar to the original COF-102, except that its linkers are formed by substituting all atoms in the boroxine (B3O3) rings with carbon atoms. This leads to enhanced thermal conductivity of Wm-1K-1, which is % larger than the reported value of pristine COF-102 ( Wm-1K-1 Liu et al. (2016)). According to the literature, low thermal conductivity of COF-102 primarily originates from vibrational mismatches between B and O atoms Liu et al. (2016). In contrast, such mismatches can be avoided in our derivatives due to the presence of all-carbon linkers.
We also calculated the temperature-dependent thermal conductivity of COF-102 derivatives for temperatures ranging from 100 to 400 K. Results are given in Fig. 2(a)–(c). In COF-102-1, thermal conductivity decreases monotonically with increasing temperature, exhibiting a dependence over the entire temperature range considered here. This trend suggests that thermal phonons are predominantly scattered by anharmonicity, as is widely observed in many crystalline solids Chen (2005). On the other hand, COF-102-2 and COF-102-3 exhibit a weak temperature dependence that deviates from classical behavior. In particular, the trend observed in COF-102-3 more closely resembles that of amorphous materials, although calculated X-ray diffraction patterns show distinct peaks at specific angles, indicating a high degree of crystallinity (see Supporting Information Sec. SIII) Robbins and Minnich (2015).
Following Refs. Ma and Tian (2019); Ma et al. (2021); Hu et al. (2024), to identify origin of the observed behavior, we computed the probability density estimate (PDE) projected onto the plane. PDE is a measure of spatial probability of finding atoms at given spatial locations by projecting all atomic coordinates onto a chosen plane. As shown in Fig. 2(d), COF-102-1 exhibits uniform pore geometries with a high PDE (yellow). In contrast, Figs. 2(e) and 2(f) show that the PDE progressively decreases (shifting toward blue) from COF-102-2 to COF-102-3, indicating increased pore irregularity compared to COF-102-1. We note that PDE maps for bor topologies exhibit similar trends across the derivatives, as compared to ctn topologies (See Supporting Information Sec. SIV for additional PDE maps of bor structures). Prior studies have established that lower PDE values are associated with greater chain rotation, which reduces overlap of periodic atomic positions, and that such rotational disorder introduces additional vibrational scattering, leading to a weaker temperature dependence in thermal conductivity Ma and Tian (2019); Ma et al. (2021). Based on these findings, we suggest that the reduced PDE in low-density COFs could be attributed to increased linker rotations, which contribute to the anomalous temperature dependence observed in these crystalline structures.
To gain insight into microscopic details, we calculated spectral energy density (SED) using Eq. 2 to extract phonon dispersion and spectral phonon MFPs Thomas et al. (2010). Phonon dispersions over selected frequency ranges are shown in Fig. 3, while full dispersions are given in Supporting Information Sec. SV. As shown in Fig. 3(a), COF-102-1 exhibits well-defined phonon modes up to 5 THz, indicating the presence of long-lived thermal phonons. Calculated group velocities near the point of the longitudinal acoustic (LA) and transverse acoustic (TA) branches are 10 km s-1 and 6.5 km s-1, respectively. In contrast, Fig. 3(b) and 3(c) show that COF-102-2 and COF-102-3 exhibit significantly broadened phonon modes in the sub-THz region. Moreover, these structures show reduced group velocities, as shown by the softening of low-frequency acoustic phonon modes compared to COF-102-1. The observed phonon softening and band broadening in them are attributed to previously discussed increased chain rotations, as reported in prior studies Ma and Tian (2019); Ma et al. (2018). We additionally note that, particularly in COF-102-2 and COF-102-3, bor topologies generally exhibit reduced LA group velocities in the low-frequency limit by 30% compared to ctn topologies, along with smaller vibrational density of states (VDOS) peak intensity (see Supporting Information Sec. SVI). We attribute our observations to bending deformations of trigonal linkers, which introduce additional vibrational scattering. Previous studies have reported that flexible C–C bonds in carbon-based structures can induce structural deformations, enhancing phonon scattering and thereby reducing phonon lifetimes and thermal conductivity, which is consistent with our observations Zhang (2018a, b); Ying et al. (2021). It is also worth noting that speeds of sound in COF-102 derivatives depend on linker length, in contrast to MOFs. This behavior can be understood by applying a simple 1D chain model as discussed in Ref. Kamencek et al. (2019) (see Supporting Information Sec. SVII for detailed discussions).
Next, we computed the spectral MFPs following the procedure in Ref. Robbins and Minnich (2015). Figure 3(d) shows the MFP as a function of frequency for all COF-102 derivatives, computed using velocities extracted from the SED results, along with lifetimes yielded by Lorentzian line-shape fit in Eq. 3. We observe that COF-102-1 can support sub-THz vibrational modes with MFPs exceeding 1 m, resulting from their high group velocities and long lifetimes. Noting that the specific heat of COF-102-1 is larger than the others owing to its substantially large mass density, such long MFPs, along with large group velocities, imply that these modes may significantly contribute to the observed orders of magnitude increase in thermal conductivity in COF-102-1, as also reported in -conjugated 2D porous polymers Ma et al. (2018).
We further calculated the cumulative thermal conductivity using the non-equilibrium molecular dynamics (NEMD) method. Unlike the EMD method, NEMD predictions are strongly affected by characteristic length of simulated system, known as size effect. Figure 4(a) illustrates the temperature profile of COF-102-1 (ctn), along with a schematic of our simulation setup. In our NEMD simulations, briefly, two fixed regions were placed at the ends of simulation cell, and two Langevin thermostats with a 60 K temperature difference were used to generate heat flux. Finally, thermal conductivity was extracted from the best linear fit of steady-state temperature profile computed by Eq. 4 (see Sec. II Methods for details).
Figure 4(b) shows computed thermal conductivity versus , together with cumulative thermal conductivity versus MFPs. According to NEMD simulations, the thermal conductivity of COF-102-1 (ctn) increases from 8.8 Wm-1K-1 to 18.5 Wm-1K-1 as extends from nm to nm. To estimate bulk thermal conductivity from NEMD (), we extrapolated the versus the data, yielding a value of Wm-1K-1, which is in a quantitative agreement with our EMD results (see Supporting Information Sec. SVIII).
We next seek to extract thermal conductivity accumulation versus MFPs using the NEMD results by solving an inverse problem described in Ref. Wei et al. (2014). Since the solution of an inverse problem is not unique, we reconstructed all possible solutions using the approach established in prior studies Robbins et al. (2019); Kim et al. (2021). Briefly, this approach enables reconstruction of MFP distribution that satisfies smoothness criteria of priors, while reproducing simulated data by incorporating uncertainty (see Supporting Information Sec. SIX for descriptions). We obtained the probability distribution of the thermal conductivity accumulation function along with the 95% credible interval using Bayesian inference with a Metropolis–Hastings Markov chain Monte Carlo method Robbins et al. (2019); Kim et al. (2021).
The reconstructed MFP spectra are given as shaded region in Fig. 4(b). We observe that phonons with MFPs on the order of 100 nm contribute significantly to the heat conduction. A quantitative analysis based on the mean distribution of the MFP spectra indicates that % of the thermal conductivity is contributed by phonons with MFPs exceeding 100 nm. For acoustic phonons, modes with hundreds of nanometer scale MFP are supported up to THz, as evidenced by SED results. This finding indicates that acoustic phonons with frequencies below THz can substantially contribute to the heat conduction in COF-102-1.
We note that the MFP distribution of COF-102-1 with significant contribution from MFPs on the order of 100 nm is similar to that in a perfect PE crystal, despite quantitatively smaller thermal conductivity of COF-102-1. More precisely, prior ab-initio calculation on a perfect PE crystal, which reported thermal conductivity of Wm-1K-1, showed that phonon MFPs range from 1 nm to 1 m, with % of the heat conduction carried by phonon modes with MFP nm, in a quantitative agreement with our results Shulumba et al. (2017). Note that the observed contributions from MFPs nm is also quantitatively similar to those experimentally resolved in macroscopic films; % for partially ordered PE Robbins et al. (2019) and roughly % for highly ordered PE Kim et al. (2022).
We first discuss the role of chain rotation in the thermal conduction of organic materials. Prior MD simulations have demonstrated that chain rotation of phenyl rings tends to reduce thermal conductivity of polymers, as rotational disorder softens acoustic branches and shortens phonon MFPs Ma and Tian (2019). For single-chain polymers, in particular, the impact of the chain rotation on suppressing thermal conductivity was reported to be more pronounced as the chain length increases Liu and Yang (2012). In the case of bulk aromatic polymers, thermal conductivity is higher than that of their single-chain counterpart, owing to - stacking between adjacent chains, which constrains random motion of phenyl rings Yang et al. (2023). Similar to these polymeric materials, our 3D COFs with shorter phenyl-based linkers also exhibit reduced chain rotations and enhanced thermal conductivity. These results suggest that restricting rotation of phenyl rings provides a general design strategy for improving intrinsic thermal conductivity of organic solids.
Finally, we discuss potential strategies that could further enhance the thermal conductivity of COFs. Recent studies have highlighted the importance of phonon focusing accompanied by anisotropic bonding in polymer crystals for achieving the intrinsic upper limits of thermal conductivity in them Cheng et al. (2019). For instance, a perfect PT crystal has been shown to exhibit higher uniaxial thermal conductivity along the chain direction compared to silicon, despite its phonon lifetimes being an order of magnitude lower. This enhancement is attributed to exceptional phonon focusing effects, where high phonon group velocities are aligned along the chain axis. Furthermore, it was demonstrated that the thermal conductivity of PT along the chain direction would drop significantly from 197 Wm-1K-1 to 34 Wm-1K-1 if the material were to become isotropic. Supporting observations in Ref. Cheng et al. (2019), our findings suggest that the thermal conductivity of COFs has not yet reached its upper bound, as COF-102 derivatives investigated in this study are purely isotropic. We anticipate that thermally anisotropic COFs that have anisotropic bonding effect may exhibit higher thermal conductivity, which is a topic of future work.
IV Conclusions
In summary, we report the thermal transport properties of 3D COFs using MD simulations. Among COF-102 derivatives with varied linker lengths and topologies, we found that the derivatives with high mass density and elastic stiffness, COF-102-1, exhibit the thermal conductivity of Wm-1K-1, nearly a two-fold increase relative to original COF-102. We attribute the increase in the thermal conductivity to suppressed chain rotation, leading to thermal phonons primarily scattered by anharmonicity. The computed thermal conductivity accumulation functions based on the NEMD results show that heat is predominantly conducted by low-frequency phonons with MFP exceeding 100 nm. Our work provides comprehensive insights into engineering heat conduction in COFs for thermal management applications.
Supporting Information
Structural and physical properties of COF-102 derivatives, details of EMD thermal conductivity calculations, XRD patterns, additional PDE maps for bor structures, full phonon dispersion relations using SED analysis, effect of bending deformation on VDOS, dependence of speeds of sound, extrapolation of bulk thermal conductivity in NEMD simulations, MFP reconstruction using Bayesian inference.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT)(No. RS-2023-00211070) and the National Supercomputing Center with supercomputing resources including technical support (KSC-2023-CRE-0537). The authors thank Hyemin Kim for discussions.